OpenStudy (anonymous):

find the sum of the integers from 20 to 80, inclusive.

OpenStudy (anonymous):

s = (n/2)(a + l) = 61/2 (20 + 80) = 30.5 * 100 = 3050

OpenStudy (anonymous):

not to be dense, but why is it 61/2? where did the 62 come from? otherwise i understand the formula...

OpenStudy (jamesj):

Because if you count carefully there are 61 numbers from 20 to 80, not 60. For the same reason that from 10 to 20 there are not ten integers but 11.

OpenStudy (anonymous):

OOOHHHH.wow.